# Sample Complexity Analysis of Multi-Target Detection via Markovian and Hard-Core Multi-Reference Alignment

### v2

@@ 1. Introduction
where * denotes linear convolution, $q$ is the number of signal occurrences, and $\varepsilon\sim\mathcal{N}(0,\sigma^{2}I_{LM})$ is additive Gaussian noise. Each unknown location $t_{i}\in\{0,1,\ldots,LM-1\}$ is represented by a one-hot vector $S(t_{i})\in\{0,1\}^{LM}$, which shifts $X$ so that its first entry is aligned at position $t_{i}$ in $Z$. While both the signal $X$ and the set of positions $\{t_{i}\}_{i=1}^{q}$ are unknown, we treat the positions as nuisance parameters and focus on estimating $X$. Throughout the introduction, we describe the one-dimensional MTD model with $X\in\mathbb{R}^{L}$ and measurement $Z\in\mathbb{R}^{LM}$ for clarity of exposition. However, our main results also apply in a two-dimensional setting, where the signal is $X\in\mathbb{R}^{L\times L}$ and the measurement is $Z\in\mathbb{R}^{ML\times ML}$. This extension is introduced in Section 5.2.
@@ 1.2. Motivating application: Cryo-electron microscopy
The MTD framework was recently proposed as a computational approach for recovering small structures that lie beyond the capabilities of current methodologies. Although heuristic algorithms have been developed, most notably those based on the method of moments and approximate expectation–maximization [13, 29, 12], the theoretical and statistical foundations of signal recovery in this regime remain largely unexplored [7]. To be more precise, some statistical properties of cryo-EM have been analyzed in [9, 16, 15], under the assumption that particle locations are known. However, the full problem with unknown locations has not been addressed, and the present work constitutes a first step toward this more general and realistic formulation.
@@ 1.3.1. Description of each patch
where $\varepsilon_{k}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}\mathcal{N}(0,\sigma^{2}I_{L})$, $\xi\in\{0,\dots,L-1\}$. Here, $\tilde{X}\left[\xi-i_{k}^{(1)}\bmod 2L\right]$ corresponds to the prefix entries of the first (possible) instance of $X$ starting in the patch $Z_{k}$, while $\tilde{X}\left[\xi+L-i_{k}^{(0)}\right]$ corresponds to the suffix entries of the (possible) cropped instance of $X$ that started in patch $Z_{k-1}$. The indices $\{(i_{k}^{(0)},i_{k}^{(1)})\}_{k=1}^{M}$ are random variables determined by the unknown placements of $X$ in $Z$, and they introduce statistical dependencies between patches.
@@
where $g_{k}$ is drawn i.i.d. from a distribution $\pi$ over a compact group $G$ acting on $\mathbb{R}^{L}$. A central insight from the MRA literature is that if the orbit of $X$ is uniquely determined by its moments up to order $n_{\min}$—the moment order cutoff—then reliable recovery in the low-SNR regime requires
@@ 1.3.3. Main assumptions and relation to prior work
All prior analyses of the MTD model have focused on the well-separated regime, in which signal occurrences are assumed to be separated by at least $L$ coordinates [12, 7]. This assumption greatly simplifies the problem, and under this setting, it has been shown that the sample complexity of the MTD model can be upper bounded by $\omega(\sigma^{6})$, via explicit recovery from third-order moments [12]. However, it remains unclear whether such guarantees continue to hold in more general settings that relax the separation condition, namely, when signal occurrences do not overlap but are otherwise arbitrarily positioned.
In this work, we focus on the more challenging non-overlapping regime, where signal instances may appear arbitrarily close to each other, subject only to the condition that they do not overlap. For the one-dimensional MTD problem, we model the distances between consecutive signal occurrences as independent draws from a Geometric distribution with parameter $\lambda\in(0,1)$. This assumption is natural for modeling sparsely distributed particles and is essential for endowing the resulting process with a Markov structure. We emphasize that key parts of our analysis, most notably Theorem 3.4, remain valid under weaker conditions, requiring only that patches sufficiently far apart exert limited statistical influence on each other (see Definition 5.5). A natural extension to two dimensions is obtained by assuming a hard-core spatial model, in which the same principle of weak long-range correlations ensures that distant patches exert only limited statistical influence on each other (Sections 56).
@@ 1.4. Main contributions
These results imply that, under the stated assumptions, the sample complexity of the MTD model is no greater than that of its associated i.i.d. MRA model. In particular, if orbit recovery for the associated i.i.d. MRA model is possible from $n_{\min}$-th order moments with sample complexity $\omega(\sigma^{2n_{\min}})$, then the same bound holds for the MTD model.
@@ 1.5. Work structure
The paper is organized as follows. Section 2 introduces the method of moments together with the feature-map framework and estimator definitions that we use throughout. Section 3 develops the one-dimensional MTD model: we formalize the measurement model, split it into patches, and show that the induced sequence of patches forms a Markovian MRA model (Theorem 3.4). Section 4 proves that the Markovian MRA model converges to the classical i.i.d. MRA model under exponential mixing time and derives the resulting sample-complexity guarantees; we also state the implications of this result for the method of moments (Corollary 4.3). Section 5 turns to two dimensions, introducing the hard-core placement model, admissibility on the grid graph, and the exponential mixing property for the induced random field. Section 6 establishes the corresponding 2D results: an induced (non-i.i.d.) MRA representation with exponential mixing (Theorem 6.2) and convergence to the i.i.d. MRA setting with matching sample complexity up to logarithmic factors (Theorem 6.3). Finally, Section 7 discusses implications, open questions, and extensions, including algebraic-structure variants and connections to cryo-EM and cryo-ET. All proofs are deferred to the appendices.
@@ 2. Preliminaries
The goal of this work is to characterize the sample complexity of the MTD model, defined as the number of patches $M$ needed to reliably recover the signal $X$ in the low-SNR regime. Although the measurement contains $q$ occurrences of $X$, we express complexity in terms of $M$. Under our asymptotic framework, the distribution of inter-signal spacings—and hence the stationary law of the induced Markov chain—remains fixed. Equivalently, the density parameter $\lambda$, which specifies the expected number of signal occurrences per patch, is treated as constant as $M\to\infty$. Thus, $q\asymp\lambda M$, so asymptotics in $M$ translate directly to asymptotics in $q$. The stationary distribution determined by $\lambda$ governs the typical patch structure and thereby indirectly shapes the sample complexity.
@@ 2.1. Method of moments
    Noiseless group-transformed moments. For a compact group $G$ acting on $\mathbb{R}^{L}$, let $g\sim\pi$ be a random element drawn from a distribution $\pi$ over $G$, and let $P$ be a fixed projection operator. The order-$n$ moment of the noiseless group-transformed signal is
    (5) $M_{X,\pi}^{(n)}\;\triangleq\;\mathbb{E}_{g\sim\pi}\!\left[\big(P(g\cdot X)\big)^{\otimes n}\right],$
@@
    $\big(M_{X,\pi}^{(n)}\big)_{i_{1},\ldots,i_{n}}\;=\;\mathbb{E}_{g\sim\pi}\!\left[\big(P(g\cdot X)\big)_{i_{1}}\cdots\big(P(g\cdot X)\big)_{i_{n}}\right].$
@@
    $M_{Y}^{(n)}\;\triangleq\;\mathbb{E}[Y^{\otimes n}]\;=\;\mathbb{E}_{g\sim\pi,\,\varepsilon}\!\left[\big(P(g\cdot X)+\varepsilon\big)^{\otimes n}\right],$
@@
    $\widehat{M}^{(n)}\;\triangleq\;\frac{1}{M}\sum_{k=1}^{M}(Y_{k})^{\otimes n},$
    serving as an estimator of $M_{Y}^{(n)}$.
A fundamental result in the MRA literature is that if the smallest integer $n$ such that $M_{X,\pi}^{(n)}$ uniquely determines the orbit of $X$ is $n_{\min}$, then the sample complexity in the low-SNR regime scales as
@@
Thus, $n_{\min}$ serves as the key statistical threshold controlling the asymptotic difficulty of orbit recovery.
@@ 2.2. Feature maps
which allows recovery of the noiseless moment $M_{X,\pi}^{(n)}$. On the other hand, if the objective is to directly reconstruct $X$ itself, then the natural choice of feature map is the identity transformation $\mathcal{F}(X)\triangleq X$. This general framework allows us to express our results in a way that applies both to moment-based recovery and to direct signal estimation, thereby unifying the treatment of these problems under a common formalism.
@@ 3.1.1. Model formulation
where, for each $i\in\{1,\ldots,q\}$, the location of the $i$-th occurrence of the signal $X$ in the measurement is determined by the parameter $t_{i}\in\{0,\ldots,L(M-1)\}$. The linear convolution of the one-hot vector $S(t_{i})$ with $X$ is given by
@@ 3.3.1. The stationary distribution of the Markov chain
Theorem 3.4 highlights a trade-off governed by the density parameter $\lambda$, as seen in the stationary distribution (16). When $\lambda$ is small, patches frequently contain no signal, so the average number of signal occurrences per patch, equivalently, the signal density, is low. This scarcity typically dominates the MSE when the number of patches $M$ is fixed. On the other hand, small $\lambda$ also implies that signals are far apart, which enhances mixing: most patches are empty, so consecutive patches are nearly independent. Note, however, that a larger $\lambda$ is generally preferable despite slowing the mixing. The recovery of the underlying signal $X$ relies on the stationary distribution $\pi$, and this recovery improves with higher signal density. By contrast, estimation of $\pi$ itself from the data, which is indeed easier when $\lambda$ is small, is not the limiting factor from an information-theoretic perspective. We emphasize, however, that, in practical applications, $\lambda$ is typically dictated by the data generation process and is not under direct control.
@@
Recall that following [2], the convergence rate $a(\sigma)$ is typically a polynomial function of the noise level $\sigma$, and is given by $a(\sigma)=\sigma^{2n_{\min}}$, where $n_{\min}$ is the order of the lowest moment that determines the orbit of the signal under the group action.
@@ 4.2. Implication to the method of moments and the MTD sample complexity
(17) $\widehat{M}_{\mathrm{MRA}}^{n}(\{Y_{k}\}_{k=1}^{M}):=\frac{1}{M}\sum_{k=1}^{M}(Y_{k})^{\otimes n}.$
@@
(18) $\widehat{M}_{\mathrm{MTD}}^{n}(\{Z_{k}\}_{k=1}^{M}):=\frac{1}{M}\sum_{k=1}^{M}(Z_{k})^{\otimes n}.$
@@
Let $n\in\mathbb{N}$, and denote by $M^{n}_{\mathrm{MRA}}=\mathbb{E}[Y^{\otimes n}]$ the $n$-th order moment of the i.i.d. MRA model, where $Y$ is as defined in (3). Let $\widehat{M}_{\mathrm{MTD}}^{n}$ be the empirical $n$-th order moment computed from $M$ consecutive patches $\{Z_{k}\}_{k=1}^{M}$ extracted from the MTD model, as defined in (18). Assume that the conditions of Theorem 4.2(ii) hold, and that the moment map $(\cdot)^{\otimes n}:\mathbb{R}^{L}\to\mathbb{R}^{L^{\otimes n}}$ is bounded on the support of the observations.
@@
    $\limsup_{M\to\infty}\frac{\mathbb{E}\left[\left\|\widehat{M}_{\mathrm{MTD}}^{n}(\{Z_{k}\}_{k=1}^{M})-M^{n}_{\mathrm{MRA}}\right\|^{2}\right]}{a(\sigma)/M}\leq C,$
    where $a(\sigma)$ is the convergence rate from Definition 4.1, and $C$ depends on $n$, $\|(M^{n}_{\mathrm{MRA}})^{\otimes n}\|_{\infty}$, and the mixing properties of the underlying Markov chain.
@@
    Furthermore, suppose that $n_{\min}$ is the minimal moment order required for orbit recovery in the i.i.d. MRA model (14), in the sense that the orbit of $\overline{X}$ under the group action is uniquely determined by $\{M^{n}_{\mathrm{MRA}}\}_{n=0}^{n_{\min}}$. Then, in the low-SNR regime $\sigma^{2}\to\infty$, the sample complexity of the method of moments applied to MTD data satisfies:
@@
Corollary 4.3 shows that the method of moments retains its asymptotic optimality when applied to dependent MTD data. Specifically, the empirical moment estimator $\widehat{M}^{n}_{\mathrm{MTD}}$ converges to the true i.i.d. moment $M^{n}_{\mathrm{MRA}}=\mathbb{E}[Y^{\otimes n}]$ with the same convergence rate as in the i.i.d. MRA model. This follows directly from Theorem 4.2(ii), which guarantees that empirical average estimators behave similarly under Markovian and i.i.d. sampling. The practical consequence is significant: for signals identifiable from moments up to order $n$, it suffices to estimate those $n$ moments using the MTD data. Thus, the method of moments applied to the MTD patches achieves the same sample complexity scaling as in the i.i.d. model, namely, $M=\omega(\sigma^{2n_{\min}})$ in the low-SNR regime. We stress that in this section we focus on the number of samples required to recover the moments from empirical data. Previous results (see [2, 14, 9]) show how to recover the signal $X$ itself from sufficiently accurate estimators of the moments, and therefore our sample complexity results apply to estimation of $X$ itself by plugin.

### v3

@@ 1. Introduction
where * denotes linear convolution, $q$ is the number of signal occurrences, and $\varepsilon\sim\mathcal{N}(0,\sigma^{2}I_{LM})$ is additive Gaussian noise. Each unknown location $t_{i}\in\{0,1,\ldots,L(M-1)\}$ is represented by a one-hot vector $S(t_{i})\in\{0,1\}^{L(M-1)+1}$, which shifts $X$ so that its first entry is aligned at position $t_{i}$ in $Z$. While both the signal $X$ and the set of positions $\{t_{i}\}_{i=1}^{q}$ are unknown, we treat the positions as nuisance parameters and focus on estimating $X$. Throughout the introduction, we describe the one-dimensional MTD model with $X\in\mathbb{R}^{L}$ and measurement $Z\in\mathbb{R}^{LM}$ for clarity of exposition. However, our main results also apply in a two-dimensional setting, where the signal is $X\in\mathbb{R}^{L\times L}$ and the measurement is $Z\in\mathbb{R}^{ML\times ML}$. This extension is introduced in Section 5.
@@ 1.2. Motivating application: Cryo-electron microscopy
The MTD framework was recently proposed as a computational approach for recovering small structures that lie beyond the capabilities of current methodologies. Although heuristic algorithms have been developed, most notably those based on the method of moments and approximate expectation–maximization [14, 31, 32, 13], the theoretical and statistical foundations of signal recovery in this regime remain largely unexplored [8, 6]. To be more precise, some statistical properties of cryo-EM have been analyzed in [10, 17, 16], under the assumption that particle locations are known. However, the full problem with unknown locations has not been addressed, and the present work constitutes a first step toward this more general and realistic formulation.
@@ 1.3.1. Description of each patch
where $\varepsilon_{k}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}\mathcal{N}(0,\sigma^{2}I_{L})$, $\xi\in\{0,\dots,L-1\}$. Here, $\tilde{X}\left[\xi-i_{k}^{(1)}\bmod 2L\right]$ corresponds to the prefix entries of the first (possible) instance of $X$ starting in the patch $Z_{k}$, while $\tilde{X}\left[\xi+L-i_{k}^{(0)}\right]$ corresponds to the suffix entries of the (possible) cropped instance of $X$ that started in patch $Z_{k-1}$. The indices $\{(i_{k}^{(0)},i_{k}^{(1)})\}_{k=1}^{M}$ are random variables determined by the unknown placements of $X$ in $Z$, and they introduce statistical dependencies between patches. See Figure 2 for illustration of the patch construction.
Figure 2. Illustration of the patch construction in (3). Starting from the ground-truth signal $X$, we form its zero-padded version $\tilde{X}=[X,0_{L}]\in\mathbb{R}^{2L}$ (panel (a)). Two copies of $\tilde{X}$ are then cyclically shifted (panel (b)), corresponding to the two possible signal contributions to a patch: a suffix of a signal instance that started in the previous patch, and a prefix of a signal instance that starts in the current patch. Applying the projection operators $P_{0}$ and $P_{1}$ extracts the relevant $L$-dimensional portions of these shifted padded signals (panel (c)). Their sum yields the noiseless patch of the observed Markovian patch $Z_{k}$ (panel (d)).
@@
where $g_{k}$ is drawn i.i.d. from a distribution $\pi$ over a compact group $G$ and $P$ is a linear projection operator. More broadly, this model belongs to the general framework of orbit recovery, in which one seeks to estimate an unknown signal from noisy observations of transformed copies of that signal under a group action. Classical examples include MRA, where the group consists of cyclic shifts, and cryo-EM, where the group is the rotation group. In this language, the patching construction used here reduces the original MTD problem to an orbit-recovery problem associated with a suitable group action on a zero-padded signal. Indeed, each patch is expressed as a combination of projected transformed copies of the padded signal, placing the estimation problem within the broader orbit-recovery paradigm.
A central insight from the orbit recovery and MRA literature is that if the orbit of $X$ is uniquely determined by its moments up to order $n_{\min}$—the moment order cutoff—then reliable recovery in the low-SNR regime requires
@@ 1.3.3. Main assumptions and relation to prior work
All prior analyses of the MTD model have focused on the well-separated regime, in which signal occurrences are assumed to be separated by at least $L$ coordinates [13, 8, 6]. This assumption greatly simplifies the problem, and under this setting, it has been shown that the sample complexity of the MTD model can be upper bounded by $\omega(\sigma^{6})$, via explicit recovery from third-order moments [13]. However, it remains unclear whether such guarantees continue to hold in more general settings that relax the separation condition, namely, when signal occurrences do not overlap but are otherwise arbitrarily positioned.
In this work, we focus on the more challenging non-overlapping regime, where signal instances may appear arbitrarily close to each other, subject only to the condition that they do not overlap. For the one-dimensional MTD problem, we model the distances between consecutive signal occurrences as independent draws from a Geometric distribution with parameter $\lambda\in(0,1)$. This assumption is natural for modeling sparsely distributed particles and endows the resulting process with a Markov structure. We emphasize that key parts of our analysis, most notably Theorems 3.4 and 4.2, remain valid under different assumptions on the particle location distribution, requiring only that patches sufficiently far apart exert limited statistical influence on each other (see Definition 5.5). A natural extension to two dimensions is obtained by assuming a hard-core spatial model, in which the same principle of weak long-range correlations ensures that distant patches exert only limited statistical influence on each other (Sections 56).
@@ 1.4. Main contributions
These results imply that, under the stated assumptions, the sample complexity of the MTD model is no greater than that of its associated i.i.d. MRA model. In particular, if orbit recovery for the associated i.i.d. MRA model is possible from $n_{\min}$-th order moments with sample complexity $\omega(\sigma^{2n_{\min}})$, then the same bound holds for the MTD model. In Section 7 we provide empirical evidence that moments up to order $3$ suffice for recovery in the induced i.i.d. MRA model, supporting $n_{\min}=3$. Hence, when particle location is infeasible (e.g. in the low-SNR regime), recovery from MTD measurements requires $\omega(\sigma^{6})$ patches.
@@ 1.5. Work structure
The paper is organized as follows. Section 2 introduces the method of moments together with the feature-map framework and estimator definitions that we use throughout. Section 3 develops the one-dimensional MTD model: we formalize the measurement model, split it into patches, and show that the induced sequence of patches forms a Markovian MRA model (Theorem 3.4). Section 4 proves that the Markovian MRA model converges to the classical i.i.d. MRA model under exponential mixing time and derives the resulting sample-complexity guarantees; we also state the implications of this result for the method of moments (Corollary 4.3). Section 5 turns to two dimensions, introducing the hard-core placement model, admissibility on the grid graph, and the exponential mixing property for the induced random field. Section 6 establishes the corresponding 2D results: an induced (non-i.i.d.) MRA representation with exponential mixing (Theorem 6.2) and convergence to the i.i.d. MRA setting with matching sample complexity up to logarithmic factors (Theorem 6.3). Section 7 presents empirical simulations that (i) demonstrate recovery in the induced i.i.d. MRA model from second- and third-order moments and (ii) illustrate the convergence of Markovian patch moments to their i.i.d. counterparts, supporting the theoretical scaling laws. Finally, Section 8 discusses implications, open questions, and extensions, including algebraic-structure variants and connections to cryo-EM and cryo-ET. All proofs are deferred to the appendices.
@@ 2. Preliminaries
The goal of this work is to characterize the sample complexity of the MTD model, defined as the number of patches $M$ needed to reliably recover the signal $X$ in the low-SNR regime. Although the measurement contains $q$ occurrences of $X$, we express complexity in terms of $M$. Under our asymptotic framework, the distribution of inter-signal spacings—and hence the stationary law of the induced Markov chain—remains fixed. Equivalently, the density parameter $\lambda$, which specifies the expected number of signal occurrences per patch, is treated as constant as $M\to\infty$. Thus, $q\asymp\lambda M$, so asymptotics in $M$ translate directly to asymptotics in $q$. The stationary distribution determined by $\lambda$ governs the typical patch structure and thereby shapes the sample complexity.
@@ 2.1. Method of moments
    Noiseless group-transformed moments. For a compact group $G$ acting on $\mathbb{R}^{L}$, let $g\sim\pi$ be a random element drawn from a distribution $\pi$ over $G$, and let $P$ be a fixed projection operator. The $n$-order moment of the noiseless group-transformed signal is
    (5) $T_{X,\pi}^{(n)}\;\triangleq\;\mathbb{E}_{g\sim\pi}\!\left[\big(P(g\cdot X)\big)^{\otimes n}\right],$
@@
    $\big(T_{X,\pi}^{(n)}\big)_{i_{1},\ldots,i_{n}}\;=\;\mathbb{E}_{g\sim\pi}\!\left[\big(P(g\cdot X)\big)_{i_{1}}\cdots\big(P(g\cdot X)\big)_{i_{n}}\right].$
@@
    (6) $T_{Y}^{(n)}\;\triangleq\;\mathbb{E}[Y^{\otimes n}]\;=\;\mathbb{E}_{g\sim\pi,\,\varepsilon}\!\left[\big(P(g\cdot X)+\varepsilon\big)^{\otimes n}\right],$
@@
    $\widehat{T}^{(n)}\;\triangleq\;\frac{1}{M}\sum_{k=1}^{M}(Y_{k})^{\otimes n},$
    serving as an estimator of $T_{Y}^{(n)}$.
A fundamental result in the MRA literature is that if the smallest integer $n$ such that $T_{X,\pi}^{(n)}$ uniquely determines the orbit of $X$ is $n_{\min}$, then the sample complexity in the low-SNR regime scales as
@@
Thus, $n_{\min}$ serves as the key statistical parameter controlling the asymptotic difficulty of orbit recovery.
@@ 2.2. Feature maps
which allows recovery of the noiseless moment $T_{X,\pi}^{(n)}$. On the other hand, if the objective is to directly reconstruct $X$ itself, then the natural choice of feature map is the identity transformation $\mathcal{F}(X)\triangleq X$. This general framework allows us to express our results in a way that applies both to moment-based recovery and to direct signal estimation, thereby unifying the treatment of these problems under a common formalism.
@@ 3.1. The 1-D MTD model
In this subsection, we introduce the one-dimensional MTD model and the stochastic assumptions governing the placement of signal occurrences within the measurement.
@@ 3.1.1. Model formulation
where, for each $i\in\{1,\ldots,q\}$, the location of the $i$-th occurrence of the signal $X$ in the measurement is determined by the parameter $t_{i}\in\{0,\ldots,L(M-1)\}$. The linear convolution of the one-hot vector $S(t_{i})$ with $X$ is given by the vector $S(t_{i})\ast X\in\mathbb{R}^{LM}$ defined, for any $\xi\in\{0,\ldots,LM-1\}$, as
@@ 3.1.2. The distribution of signal appearances
We emphasize that our convergence results are not specific to Assumption 3.1. Indeed, any assumption on the distribution of occurrences satisfying the exponential mixing property (see Definition 5.5) would provide the same convergence guarantees. For instance, we could have chosen a 1D analogue of the hard-core model presented in Section 5.2. However, we chose to consider Assumption 3.1 for clarity and presentation purposes, since it induces a simple Markov process, and explicit computations of the convergence rate and the associated stationary distribution are possible. In addition, we note that the patching construction does not require exact knowledge of the signal length: it is enough to choose $L$ as an upper bound on the signal support. In that case, however, the non-overlapping requirement is enforced relative to this upper bound, and is therefore potentially more restrictive than if the exact support size were known.
@@ 3.2. Splitting the measurement into patches
We point out that, in view of this construction and the non-overlapping assumption, not all the group elements $g_{k}=(g_{k}^{(1)},g_{k}^{(2)})\in\mathbb{Z}^{2L}\times\mathbb{Z}^{2L}$ are admissible. In our construction, ignoring the noise variable $\varepsilon_{k}$, the patches $Z_{k}$ are modeled by
[…]
@@ 3.3.1. The stationary distribution of the Markov chain
Theorem 3.4 together with the stationary distribution (17) reveals a trade-off governed by the density parameter $\lambda$. Our reduction of the MTD problem to a Markovian MRA model relies on two key ingredients: first, the convergence of the dependent patch process to its associated stationary i.i.d. model; and second, the ability to recover the signal from that stationary model. On the one hand, as follows from the mixing estimate (16), smaller values of $\lambda$ lead to faster convergence to stationarity; see Figure 4 for an empirical illustration. On the other hand, $\lambda$ also affects the sample complexity of the stationary problem itself. Indeed, by [2, Theorem II.2] and the form of the stationary distribution $\pi$ in (17), when $\lambda$ is small, patches contain signal less frequently, so the induced random variable $Y=P(g\cdot\overline{X})$, with $g\sim\pi$, has weaker low-order moments. As a result, more patches are needed to achieve a given estimation accuracy. Thus, although smaller $\lambda$ improves the mixing properties of the underlying Markov chain, larger $\lambda$ is generally more favorable for signal recovery because it increases the signal density in the stationary model. In other words, the main information-theoretic bottleneck is not estimating the stationary distribution $\pi$, but recovering $X$ from the induced stationary observations. We emphasize, however, that in practical applications $\lambda$ is typically determined by the data-generation mechanism and is not under direct control.
@@
Recall that, following [2], the convergence rate $a(\sigma)$ is polynomial in the noise level $\sigma$. In standard orbit-recovery models, the optimal estimation rate is of the form $a(\sigma)=\sigma^{2n_{\min}}$, where $n_{\min}$ is the order of the lowest moment that determines the orbit of the signal under the group action.
@@ 4.2. Implication to the method of moments and the MTD sample complexity
(18) $\widehat{T}_{\mathrm{MRA}}^{(n)}(\{Y_{k}\}_{k=1}^{M}):=\frac{1}{M}\sum_{k=1}^{M}(Y_{k})^{\otimes n}.$
@@
(19) $\widehat{T}_{\mathrm{MTD}}^{(n)}(\{Z_{k}\}_{k=1}^{M}):=\frac{1}{M}\sum_{k=1}^{M}(Z_{k})^{\otimes n}.$
@@
Let $n\in\mathbb{N}$, and denote by $T^{(n)}_{\mathrm{MRA}}=\mathbb{E}[Y^{\otimes n}]$ the $n$-th order population moment of the i.i.d. MRA model, where $Y$ is as defined in (3). Let $\widehat{T}_{\mathrm{MTD}}^{(n)}$ be the empirical $n$-th order moment computed from $M$ consecutive patches $\{Z_{k}\}_{k=1}^{M}$ extracted from the MTD model, as defined in (19). Assume that the conditions of Theorem 4.2(ii) hold, and that the moment map $(\cdot)^{\otimes n}:\mathbb{R}^{L}\to\mathbb{R}^{L^{\otimes n}}$ is bounded on the support of the observations.
@@
    $\limsup_{M\to\infty}\frac{\mathbb{E}\left[\left\|\widehat{T}_{\mathrm{MTD}}^{(n)}(\{Z_{k}\}_{k=1}^{M})-T^{(n)}_{\mathrm{MRA}}\right\|^{2}\right]}{a(\sigma)/M}\leq C,$
    where $a(\sigma)$ is the convergence rate from Definition 4.1, and $C$ depends on $n$, $\|(T^{(n)}_{\mathrm{MRA}})^{\otimes n}\|_{\infty}$, and the mixing properties of the underlying Markov chain.
@@
    Furthermore, suppose that $n_{\min}$ is the minimal moment order required for orbit recovery in the i.i.d. MRA model (15), in the sense that the orbit of $\overline{X}$ under the group action is uniquely determined by $\{T^{(n)}_{\mathrm{MRA}}\}_{n=0}^{n_{\min}}$. Then, in the low-SNR regime $\sigma^{2}\to\infty$, the sample complexity of the method of moments applied to MTD data satisfies:
@@
Corollary 4.3 shows that the method of moments retains its asymptotic optimality when applied to dependent MTD data. Specifically, the empirical moment estimator $\widehat{T}^{(n)}_{\mathrm{MTD}}$ converges to the true i.i.d. moment $T^{(n)}_{\mathrm{MRA}}=\mathbb{E}[Y^{\otimes n}]$ with the same convergence rate as in the i.i.d. MRA model. This follows directly from Theorem 4.2(ii), which guarantees that empirical average estimators behave similarly under Markovian and i.i.d. sampling. The practical consequence is significant: for signals identifiable from moments up to order $n$, it suffices to estimate those $n$ moments using the MTD data. Thus, the method of moments applied to the MTD patches achieves the same sample complexity scaling as in the i.i.d. model, namely, $M=\omega(\sigma^{2n_{\min}})$ in the low-SNR regime. We stress that in this section we focus on the number of samples required to recover the moments from empirical data. Previous results (see [2, 15, 10]) show how to recover the signal $X$ itself from sufficiently accurate estimators of the moments, and therefore our sample complexity results apply to estimation of $X$ itself by plugin.